# Aladin Djuhera1, Amin Seffo1, Vlad C. Andrei1, Holger Boche1, and Walid Saad2

### v2

@@ Abstract
Path planning under wireless performance constraints is a complex challenge in robot navigation. However, naively incorporating such constraints into classical planning algorithms often incurs prohibitive search costs. In this paper, we propose SCoTT, a wireless-aware path planning framework that leverages vision-language models (VLMs) to co-optimize average path gains and trajectory length using wireless heatmap images and ray-tracing data from a digital twin (DT). At the core of our framework is Strategic Chain-of-Thought Tasking (SCoTT), a novel prompting paradigm that decomposes the exhaustive search problem into structured subtasks, each solved via chain-of-thought prompting. To establish strong baselines, we compare classical A* and wireless-aware extensions of it, and derive DP-WA*, an optimal, iterative dynamic programming algorithm that incorporates all path gains and distance metrics from the DT, but at significant computational cost. In extensive experiments, we show that SCoTT achieves path gains within 2% of DP-WA* while consistently generating shorter trajectories. Moreover, SCoTT’s intermediate outputs can be used to accelerate DP-WA* by reducing its search space, saving up to 62% in execution time. We validate our framework using four VLMs, demonstrating effectiveness across both large and small models, thus making it applicable to a wide range of compact models at low inference cost. We also show the practical viability of our approach by deploying SCoTT as a ROS node within Gazebo simulations. Finally, we discuss data-acquisition pipelines, compute requirements, and deployment considerations for VLMs in 6G-enabled DTs, underscoring the potential of natural language interfaces for wireless-aware navigation in real-world applications.
@@ I Introduction and Motivation
Future 6G networks will underpin a host of emerging technologies, including artificial general intelligence (AGI) [1], autonomous transportation [2], and the metaverse [3]. At the same time, 6G-enabled digital twins (DTs) will be pivotal for smart manufacturing, intelligent planning, and autonomous robotics [4]. DTs act as high-fidelity replicas of real physical systems and processes, thus encompassing not only their physical geometry but also their associated algorithms, communication protocols, and compute resources [4]. Embedding wireless ray-tracing data into DTs will become critical for capturing realistic network-layer behavior and enabling connectivity-aware decision making. In our prior work [5], for example, we developed a ray-tracing-enabled DT framework for integrated sensing and communication (ISAC)-enabled robotics that fuses high-resolution wireless data with 3D environment models. This allows for real-time, wireless-aware interactions between robots and their digital environments.
@@
This example illustrates an inherently complex problem: Path planning with side constraints beyond simple distance minimization. In general, path planning is an exhaustive search problem and traditional algorithms become inefficient when operating over large datasets or incorporating complex side constraints. This motivates research on whether path planning can be automated via LLMs. For example, S2RCQL [7] introduces a curriculum-learning approach with spatial-to-relational prompting, training the LLM to imitate reinforcement learning strategies. However, it struggles with scalability and often hallucinates in complex environments. Likewise, LLM-A* [8] augments the classical A* algorithm [9] by guiding the search with LLM-generated waypoints via few-shot prompting, reducing unnecessary exploration and increasing efficiency.
However, reducing redundancy in wireless measurement data, such as path gains, may not be straightforward, posing a problem to data-efficient LLM-based path planning, and requiring new approaches to how wireless data is represented, interpreted, and incorporated into the LLM prompt. In this work, we address this gap by introducing Strategic Chain-of-Thought Tasking (SCoTT), a novel prompting paradigm that decomposes the wireless-aware path planning problem into structured subproblems, each solved through strategic chain-of-thought (SCoT) prompting. Specifically, we employ vision-language models (VLMs) and leverage retrieval-augmented generation (RAG) [10] pipelines to process multi-modal wireless ray-tracing data (wireless heatmap images and path gain data), generated by a real-world DT implementation from [5]. To demonstrate effectiveness, we apply SCoTT across four VLMs of varying sizes (Llama-4-Scout-17B [11], Llama-3.2-11B-Vision [12], SmolVLM [13], and Granite-Vision-3.2-2B [14]) and compare against wireless-aware extensions of the classical A* algorithm. In summary, our key contributions are:
@@
  • To ensure real-world viability, we detail our ROS-based integration of SCoTT within Gazebo [15] and present navigation outcomes in scenarios involving wireless-aware path planning, obstacle avoidance, and motion control.
@@
The remainder of the paper is organized as follows: Sec. II introduces wireless-aware extensions of A*. Sec. III presents our SCoTT framework. Simulation results are shown in Sec. IV, followed by our practical ROS integration in Sec. V, and conclusions in Sec. VI.
@@ II Wireless-Aware Algorithmic Path Planning
II-A N-WA*: Naive Wireless-Aware A*.
@@
This approach is naive because even though it prefers nodes with higher path gains, it does not explicitly enforce any power thresholds, making it a heuristic-based adaptation rather than a strict constraint optimization. Thus, N-WA* inherently biases toward the shortest path. This may lead to suboptimal paths when a high-gain route is longer, but ultimately preferable. Additionally, it cannot explicitly handle scenarios where an average path gain threshold $G$ must be maintained along the entire path. However, the additional storage of path gain values does not fundamentally alter the asymptotic complexity of the algorithm such that it has identical complexity as A*, i.e. $\mathcal{O}(b^{d})$, where $b$ is the average number of neighbors per node and $d$ is the number of steps in the optimal path. This makes N-WA* still inefficient for large data points, but avoids the substantial time and memory overhead of dynamic-programming (DP) solutions, as shown in the next section.
@@
II-B DP-WA*: Dynamic Programming Wireless-Aware A*.
@@
where $g_{\max}=1$ is the maximum path gain at any position. This feasibility condition prunes states that cannot possibly meet the wireless constraint by the final time step. With that, the optimal control policy at step $k$ selects the action $u_{k}^{*}$ that minimizes the total cost-to-go from $p_{k}$ under the wireless path gain constraint and is given by
@@ 3 Output: Optimal path from $p_{0}$ to $p_{T}$ satisfying $G$
5Initialize the DP table $V(p,W)\leftarrow\infty$ for all positions $p$ and accumulated gains $W$
@@ III SCoTT: Stragetic Chain-of-Thought Tasking
To overcome these limitations, SCoTT tasks the VLM via a chain of smaller, strategy-driven prompts, thereby decomposing the complex wireless-aware path planning problem into focused subproblems, each independently solved via strategic chain-of-thought (SCoT) prompting. Unlike standard CoT, SCoT first elicits a high-level strategy for each subproblem and then uses that strategy to guide a sequence of reasoning steps, turning the monolithic task into a series of discrete, strategy-driven CoT tasks. In our wireless-aware path planning example, SCoT prompting mainly involves two steps:
@@
We decompose the planning task into consecutive SCoT prompts as follows. First, the VLM analyzes the wireless heatmap to identify high‐gain corridors of constraint-compliant waypoint regions. Then it performs three subtasks: (i) coarse path selection, which generates an initial route, (ii) focus‐area definition, which selects reduced search regions around the initial route, and (iii) fine‐grained waypoint refinement, which computes the optimized, detailed path. As shown in Fig. 2, we prompt the VLM in three stages:
@@ 2. Subtask 2:
    Focus‐Area Definition. Given the coarse path, we instruct the VLM to select $N$ focus areas, each stored as a $k$d‐tree with radius <max_distance>, serving as reduced search spaces around the initial route. Users may further adjust $N$ and <max_distance> to increase the number of focus areas, effectively broadening or narrowing the fine-grained search in the next step.
@@ 3. Subtask 3:
    Fine‐Grained Refinement. Finally, for each focus area, the model incrementally generates a detailed path using RAG-supplied ray-tracing data, ensuring the wireless path gain threshold $G$ is met. Because each prompt now covers only a small region, the context window easily holds both instructions and data, allowing vision-guided waypoint selection with precise gain constraints. By iterating over these compact areas, the model avoids large-scale hallucinations and consistently respects the wireless path gain threshold $G$.
@@
SCoTT thus efficiently decomposes a single, complex path planning query into a sequence of focused CoT tasks, each scoped to a specific focus area and guided by an explicitly elicited reasoning strategy and workflow. This targeted breakdown prevents hallucinations and avoids context overflow, while producing concise and explainable outputs. We further note that the resulting ensemble of focus areas (represented as $k$d‐trees) from subtask 2 can serve as input to classical algorithms like DP‐WA* to reduce their search space. We explore this neuro-symbolic extension in Sec. IV. Furthermore, SCoTT uses the DT’s path gain data intelligently, invoking it only during the fine‐grained refinement stage, thereby balancing data richness with prompt efficiency. We demonstrate SCoTT’s effectiveness across different VLM models in the next section.
@@ IV Simulation Results
Path 1: Across the Room Path 2: Wall to Wall Path 3: Extreme Case
A* Avg. Path Gain 0.34 0.06 0.10
@@
SCoTT Avg. Path Gain 0.72 0.41 0.62
@@
SCoTT- DP-WA* Avg. Path Gain 0.73 0.43 0.66
@@
IV-A Experimental Setup.
We evaluate SCoTT across three challenging path planning examples, involving areas of both high and low wireless path gains. We conduct all simulations using the DT implementation from [5], where a 3D model of our lab at TUM was twinned into NVIDIA Omniverse [17]. Therein, wireless ray tracing data was added using Remcom WirelessInsite [18] with accurate path gains for the twinned access point, which operates at 2.4 GHz using OFDM signaling with 1024 subcarriers, each spaced by 78.125 kHz. All dimensions in the subsequent results are given in meters and path gains are normalized. Whitespaces represent obstacles such as cupboards or tables. We leave the investigation of more complex DT environments to future work. Our main interest is the validity of generated paths and corresponding performance metrics compared to classical A*, N-WA*, and DP-WA* approaches. For SCoTT, we use Llama-4-Scout-17B as our primary model and explore additional models in the further analysis. We found that using five to seven focus areas ($N$) consistently produced the best outcomes. In addition, we present ROS simulation results within the DT environment in Sec. V to complement SCoTT’s path planning analysis with real-world navigation outcomes.
IV-B Path Planning Results.
@@ IV-B1 Path 1 - Across the Room
In Fig. 3, a trajectory with average path gain threshold $G=0.7$ is planned across the room from $n_{\text{start}}=(-2.5,-1)$ to $n_{\text{goal}}=(2.5,4.5)$, traversing through areas with high, mid, and low path gains. The classical A* algorithm naturally chooses the shortest, unobstructed path with length 7.43m, represented by the purple straight line, and thus has the lowest average path gain of 0.34. Similarly, N-WA* chooses a path with length 7.77m but deviates more toward higher gain areas, obtaining an average gain of 0.46. In contrast, the optimal DP-WA* path achieves the highest path gain of 0.75 at cost of a both increased path length at 9.15m and computation time of 76.04s, i.e., nearly 19 times higher than N-WA*. SCoTT, represented by the green path, is able to achieve a similar gain of 0.72 while slightly reducing the path length to 8.64m. The VLM-generated path in particular starts moving immediately to the higher gain areas and overlaps with DP-WA* therein before shortening its path toward $n_{\text{goal}}$ in the last few meters. Moreover, SCoTT produces the full path in 56s total, where it takes 5s to process the initial prompt and 7s, 15s, and 29s for executing subtasks 1, 2, and 3 (see Fig. 2). As noted in the previous section, DP-WA* can be significantly accelerated by reducing the search space using SCoTT’s focus areas. This results in close performance to the original DP-WA* while reducing the computation time by more than 61% to 30.01s, represented by the magenta SCoTT-DP-WA* path. This is especially useful because SCoTT can precompute focus areas and cache them for reuse in similar path-planning tasks. Thus, only SCoTT or DP paths can successfully adhere to $G$ with SCoTT achieving very close performance to DP-WA*.
@@ IV-B2 Path 2 - Wall to Wall
A more challenging scenario in a lower gain area is shown in Fig. 4, where a trajectory with threshold $G=0.4$ is requested. Intuitively, A* chooses the shortest path with length 7.07m and catastrophic gain 0.06, thus maintaining a straight line to the adjacent wall while going around the whitespace table obstacle. In contrast, all other paths, including N-WA*, take a detour into higher path gain areas to optimize for $G$. However, as before, only SCoTT and DP paths can adhere to $G$ with DP-WA* having the highest gain of 0.43, compared to SCoTT with 0.41. In this example, SCoTT-DP-WA* is identical to DP-WA* whereas for Path 1 SCoTT’s reduced focus areas may have omitted some few, but more optimal points, resulting in a slight difference compared to DP-WA*. In addition, as before, SCoTT tends to reduce the total path length while being faster than DP-WA* at around 52s. In addition, SCoTT-DP-WA* reduces DP-WA*’s computation time by more than 62% to 29.17s.

### v3

# Aladin Djuhera1, Amin Seffo1, Vlad C. Andrei1, Holger Boche1, Walid Saad2

@@ Abstract
Path planning under wireless performance constraints is a complex challenge in robot navigation. However, naively incorporating such constraints into classical planning algorithms often incurs prohibitive search costs. In this paper, we propose SCoTT, a wireless-aware path planning framework that leverages vision-language models (VLMs) to co-optimize average path gains and trajectory length using wireless heatmap images and ray tracing data from a digital twin (DT). At the core of our framework is Strategic Chain-of-Thought Tasking (SCoTT), a novel prompting paradigm that decomposes the exhaustive search problem into structured subtasks, each solved via chain-of-thought prompting. To establish strong baselines, we compare classical A* and wireless-aware extensions of it, and derive DP-WA*, an optimal, iterative dynamic programming algorithm that incorporates all path gains and distance metrics from the DT, but at significant computational cost. In extensive experiments, we show that SCoTT achieves path gains within 2% of DP-WA* while consistently generating shorter trajectories. Moreover, SCoTT’s intermediate outputs can be used to accelerate DP-WA* by reducing its search space, saving up to 62% in execution time. We validate our framework using four VLMs, demonstrating effectiveness across both large and small models, thus making it applicable to a wide range of compact models at low inference cost. We also show the practical viability of our approach by deploying SCoTT as a ROS node within Gazebo simulations. Finally, we discuss data-acquisition pipelines, compute requirements, and deployment considerations for VLMs in 6G-enabled DTs, underscoring the potential of natural language interfaces for wireless-aware navigation in real-world applications.
@@ I Introduction and Motivation
Future 6G networks will underpin a host of emerging technologies, including artificial general intelligence (AGI) [1], autonomous transportation [2], and the metaverse [3]. At the same time, 6G-enabled digital twins (DTs) will be pivotal for smart manufacturing, intelligent planning, and autonomous robotics [4]. DTs act as high-fidelity replicas of real physical systems and processes, thus encompassing not only their physical geometry but also their associated algorithms, communication protocols, and compute resources [4]. Embedding wireless ray tracing data into DTs will become critical for capturing realistic network-layer behavior and enabling connectivity-aware decision making. In our prior work [5], for example, we developed a ray tracing-enabled DT framework for integrated sensing and communication (ISAC)-enabled robotics that fuses high-resolution wireless data with 3D environment models. This allows for real-time, wireless-aware interactions between robots and their digital environments.
@@
This example illustrates an inherently complex problem: Path planning with side constraints beyond simple distance minimization. In general, path planning is an exhaustive search problem and traditional algorithms become inefficient when operating over large datasets or incorporating complex side constraints. This motivates research on whether path planning can be automated via LLMs. For example, S2RCQL [7] introduces a curriculum-learning approach with spatial-to-relational prompting, training the LLM to imitate reinforcement learning strategies. However, it struggles with scalability and often hallucinates in complex environments. Similarly, LLM-A* [8] augments the classical A* algorithm [9] by guiding the search with LLM-generated waypoints via few-shot prompting, reducing unnecessary exploration and increasing efficiency.
However, reducing redundancy in wireless measurement data, such as path gains, may not be straightforward, posing a problem for data-efficient LLM-based path planning and requiring new approaches to how wireless data is represented, interpreted, and incorporated into the LLM. In this work, we address this gap by introducing Strategic Chain-of-Thought Tasking (SCoTT), a novel prompting paradigm that decomposes the wireless-aware path planning problem into structured subproblems, each solved through strategic chain-of-thought (SCoT) reasoning [10], an extension of the popular chain-of-thought (CoT) prompting approach [11]. Specifically, we employ vision-language models (VLMs) and leverage retrieval-augmented generation (RAG) [12] to process multi-modal wireless ray tracing data (wireless heatmap images and path gain data), generated by a real-world DT implementation from [5]. To demonstrate effectiveness, we apply SCoTT across four VLMs of varying sizes (Llama-4-Scout-17B [13], Llama-3.2-11B-Vision [14], SmolVLM [15], and Granite-Vision-3.2-2B [16]) and compare against wireless-aware extensions of the classical A* algorithm. In summary, our key contributions are:
@@
  • To ensure real-world viability, we detail our robot operating system (ROS)-based integration of SCoTT within Gazebo [17] and present wireless-aware navigation outcomes in scenarios involving obstacle avoidance and motion control.
@@
The remainder of the paper is organized as follows: Sec. II introduces wireless-aware extensions of A*. Sec. III presents our SCoTT framework. Simulation results are shown in Sec. IV, followed by the practical ROS integration in Sec. V and the conclusions in Sec. VI.
@@ II Wireless-Aware Algorithmic Path Planning
II-A N-WA*: Naive Wireless-Aware A*
@@
This approach is naive because even though it prefers nodes with higher path gains, it does not explicitly enforce any thresholds, making it a heuristic-based adaptation rather than a strict constraint optimization. Thus, N-WA* inherently biases toward the shortest path. This may lead to suboptimal paths when a high-gain route is longer, but ultimately preferable. Additionally, it cannot explicitly handle scenarios where an average path gain threshold $G$ must be maintained along the entire path. However, the additional storage of path gain values does not fundamentally alter the asymptotic complexity of the algorithm such that it has identical complexity as A*, i.e. $O(b^{d})$, where $b$ is the average number of neighbors per node and $d$ is the number of steps in the optimal path. This makes N-WA* still inefficient for large data points, but avoids the substantial time and memory overhead of dynamic-programming (DP) solutions, as shown in the next section.
@@
II-B DP-WA*: Dynamic Programming Wireless-Aware A*
@@
where $g_{\max}=1$ is the maximum path gain at any position. This feasibility condition prunes states that cannot possibly meet the wireless path gain constraint by the final time step. With that, the optimal control policy at step $k$ selects the action $u_{k}^{*}$ that minimizes the total cost-to-go from $p_{k}$ under the wireless path gain constraint and is given by
@@ 3 Output: Optimal path from $p_{0}$ to $p_{T}$ satisfying $G$
5
6Initialize the DP table $V(p,W)\leftarrow\infty$ for all positions $p$ and accumulated gains $W$
@@ III SCoTT: Stragetic Chain-of-Thought Tasking
To overcome these limitations, SCoTT tasks the VLM via a chain of smaller, strategy-driven prompts, thereby decomposing the complex wireless-aware path planning problem into focused subproblems, each solved independently via strategic chain-of-thought (SCoT) prompting [10]. Unlike standard CoT [11], SCoT first elicits a high-level strategy for each subproblem and then uses that strategy to guide a sequence of reasoning steps, turning the monolithic task into a series of discrete, strategy-driven CoT tasks. In our wireless-aware path planning example, SCoT prompting involves two steps:
@@
We decompose the planning task into consecutive SCoT prompts as follows. First, the VLM analyzes the wireless heatmap to identify high‐gain corridors of constraint-compliant waypoint regions. Then it performs three subtasks: (i) coarse path selection, which generates an initial route, (ii) focus‐area definition, which selects reduced search regions around the initial route, and (iii) fine‐grained waypoint refinement, which computes the optimized, fine-grained path. As shown in Fig. 2, we prompt the VLM in three stages:
@@ 2. Subtask 2:
    Focus‐Area Definition. Given the coarse path, we instruct the VLM to select $N$ focus areas, each stored as a kkd‐tree with radius <max_distance>, serving as reduced search spaces around the initial route. Users may further adjust $N$ and <max_distance> to increase the number of focus areas, effectively broadening or narrowing the fine-grained search in the next step.
@@ 3. Subtask 3:
    Fine‐Grained Refinement. Finally, for each focus area, the model incrementally generates a detailed path using RAG-supplied ray tracing data, ensuring the wireless path gain threshold $G$ is met. Because each prompt now covers only a small region, the context window easily holds both instructions and data, allowing vision-guided waypoint selection with precise gain constraints. By iterating over these compact areas, the model avoids large-scale hallucinations and consistently respects the wireless path gain threshold $G$.
@@
SCoTT thus efficiently decomposes a single, complex path planning query into a sequence of focused CoT tasks, each scoped to a specific focus area and guided by an explicitly elicited reasoning strategy and workflow. This targeted breakdown prevents hallucinations and avoids context overflow, while producing concise and explainable outputs. We further note that the resulting ensemble of focus areas (represented as kkd‐trees) from subtask 2 can serve as input to classical algorithms like DP‐WA* to reduce their search space. We explore this neuro-symbolic extension in Sec. IV. Furthermore, SCoTT uses the DT’s path gain data intelligently, invoking it only during the fine‐grained refinement stage, thereby balancing data richness with prompt efficiency. We demonstrate SCoTT’s effectiveness across different VLM models in the next section.
@@ IV Simulation Results
Path 1:
Across the Room
Path 2:
Wall to Wall
Path 3:
Extreme Case
A*
Avg. Path Gain 0.34 0.06 0.10
@@
SCoTT
Avg. Path Gain 0.72 0.41 0.62
@@
SCoTT- DP-WA*
Avg. Path Gain 0.73 0.43 0.66
@@
IV-A Experimental Setup
We evaluate SCoTT across three challenging path planning examples, involving areas of both high and low wireless path gains. We conduct all simulations using the DT implementation from [5], where a 3D model of our lab at TUM was twinned into NVIDIA Omniverse [19]. Therein, wireless ray tracing data was added using Remcom WirelessInsite [20] with accurate path gains for the twinned access point, which operates at 2.4 GHz using OFDM signaling with 1024 subcarriers, each spaced by 78.125 kHz. All dimensions in the subsequent results are given in meters and path gains are normalized. Whitespaces represent obstacles such as cupboards or tables. We leave the investigation of more complex DT environments to future work. Our main interest is the validity of generated paths and corresponding performance metrics compared to classical A*, N-WA*, and DP-WA* approaches. For SCoTT, we use Llama-4-Scout-17B as our primary model and explore additional models in the further analysis. We found that using five to seven focus areas ($N$) consistently produced the best outcomes. In addition, in Sec. V, we present corresponding ROS simulation results to complement SCoTT’s path planning analysis with real-world navigation outcomes.
IV-B Path Planning Results
@@ IV-B1 Path 1 - Across the Room
In Fig. 3, a trajectory with average path gain threshold $G=0.7$ is planned across the room from $n_{\text{start}}=(-2.5,-1)$ to $n_{\text{goal}}=(2.5,4.5)$, traversing through areas with high, mid, and low path gains. The classical A* algorithm naturally chooses the shortest unobstructed path with length 7.43m, represented by the purple straight line, and thus has the lowest average path gain of 0.34. Similarly, N-WA* chooses a path with length 7.77m but deviates more toward higher gain areas, obtaining an average gain of 0.46. In contrast, the optimal DP-WA* path achieves the highest path gain of 0.75 at cost of a both increased path length at 9.15m and computation time of 76.04s, i.e., nearly 19 times higher than N-WA*. SCoTT, represented by the green path, is able to achieve a similar gain of 0.72 while slightly reducing the path length to 8.64m. The VLM-generated path in particular starts moving immediately to the higher gain areas and overlaps with DP-WA* therein before shortening its path toward $n_{\text{goal}}$ in the last few meters. Moreover, SCoTT produces the full path in 56s total, where it takes 5s to process the initial prompt and 7s, 15s, and 29s for executing subtasks 1, 2, and 3 (see Fig. 2). As noted in the previous section, DP-WA* can be significantly accelerated by reducing the search space using SCoTT’s focus areas. This results in close performance to the original DP-WA* while reducing the computation time by more than 61% to 30.01s, represented by the magenta SCoTT-DP-WA* path. This is especially useful because SCoTT can precompute focus areas and cache them for reuse in similar path-planning tasks. Thus, only SCoTT or DP paths can successfully adhere to $G$ with SCoTT achieving very close performance to DP-WA*.
@@ IV-B2 Path 2 - Wall to Wall
A more challenging scenario in a lower gain area is shown in Fig. 4, where a trajectory with threshold $G=0.4$ is requested. Intuitively, A* chooses the shortest path with length 7.07m and catastrophic gain 0.06, thus maintaining a straight line to the adjacent wall while going around the whitespace table obstacle. In contrast, all other paths, including N-WA*, take a detour into higher path gain areas to optimize for $G$. However, as before, only SCoTT and DP paths can adhere to $G$ with DP-WA* having the highest gain of 0.43, compared to SCoTT with 0.41. In this example, SCoTT-DP-WA* is identical to DP-WA* whereas for Path 1 SCoTT’s reduced focus areas may have omitted some few, but more optimal points, resulting in a slight difference compared to DP-WA*. As before, SCoTT tends to reduce the total path length while being faster than DP-WA* at around 52s. In addition, SCoTT-DP-WA* reduces DP-WA*’s computation time by more than 62% to 29.17s.